### v1

@@ 5 Related Work
Aligning LLMs with human values remains a challenge due to biases in training data and trade-offs between usefulness and safety Zeng et al. (2024); Ding et al. (2024). Approaches such as Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Bai et al. (2022) have been proposed to improve fairness Navigli et al. (2023); Gallegos et al. (2024), safety Zou et al. (2023) and eliminate hallucinations Zhang et al. (2023); Lin et al. (2024).
@@
Jailbreak attacks threaten the safety alignment mechanisms of LLMs, potentially leading to the generation of harmful content Carlini et al. (2023); Liu et al. (2023); Yi et al. (2024); Li et al. (2025). Our study is inspired by two key methods in black-box attacks: prompt nesting and multi-turn dialogue attacks. 1) Prompt Nesting Attack. Prompt nesting bypasses security features by nesting malicious intents in normal prompts, altering LLMs’ context. DeepInception Li et al. (2023b) exploit nested scenarios, while ReNeLLM Ding et al. (2024) rewrites prompt to jailbreak based on code completion, text continuation, or form-filling tasks. MJP Li et al. (2023a) uses multi-step approaches with contextual contamination to reduce moral constraints, prompting malicious responses. 2) Multi-turn Dialogue Attack. LLMs that are safe in isolated, single-round interactions can be gradually manipulated into generating harmful outputs through multiple rounds of interaction Russinovich et al. (2024); Zhou et al. (2024); Yang et al. (2024); Cheng et al. (2024). Multi-turn dialogue attack leverages the multi-turn nature of conversational interactions to gradually erode an LLM’s content restrictions.
@@ 6 Conclusion
In this study, we enhance the understanding of jailbreak attacks by proposing a novel approach, AdVersArial meTAphoR (AVATAR) that manipulates LLMs to generate harmful content through calibrating benign metaphors, rather than generating harmful content from scratch. AVATAR consists of Adversarial Entity Mapping (AEM) and Metaphor-Driven Thinking (MDT). Among them, AEM identifies metaphors by model crowdsourcing, thus enabling the stable acquisition of suitable metaphors before jailbreaking the target LLM. MDT induces the target LLM to analyze the metaphor for harmful content generating, by introducing the metaphor calibration mechanisms, MDT further enhances the efficiency of transforming benign content into harmful forms. Experiments have demonstrated our AVATAR is effective in generating transferable adversarial attacks on multiple advanced LLMs.
@@
AVATAR framework generates metaphors using the imagination of LLMs and conducts human-like interaction to prompt LLMs to generate harmful content without complex tricks or sophisticated jailbreak templates. However, the effectiveness of AVATAR is potentially limited by:
Imagination is limited by poor real-world experience.
LLMs lack direct real-world experience and their understanding of harmful entities is predominantly textual. While this vulnerability enables our metaphor-based attack, it also limits the attack’s ability to generate truly realistic and contextually appropriate harmful attacks. To alleviate this issue, we introduce the crowdsourcing strategy to identify the nice metaphors.
@@
The safety alignment mechanisms in existing open-source and closed-source LLMs limit their deeper analysis of harmful content for better toxic entity extraction. During the Adversarial Entity Mapping stage, the attack model may extract a mix of harmful and safe entities, e.g., ethical consideration legal consequences, from the harmful query, potentially interfering with subsequent metaphorical content generation. The effectiveness of in-context learning in mitigating this issue is limited. To alleviate this issue, we introduce the word filtering mechanism to remove safe entities during the entity extracting.

### v4

@@ 5 Related Work
Aligning LLMs with human values remains a challenge due to biases in training data and trade-offs between usefulness and safety Zeng et al. (2024); Ding et al. (2024). Approaches such as Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Bai et al. (2022) have been proposed to improve fairness Navigli et al. (2023); Gallegos et al. (2024), safety Zou et al. (2023) and eliminate hallucinations Zhang et al. (2023a); Lin et al. (2024).
@@
Jailbreak attacks threaten the safety alignment mechanisms of LLMs, potentially leading to the generation of harmful content Carlini et al. (2023); Liu et al. (2023); Yi et al. (2024); Li et al. (2025). Our study is inspired by two key methods in black-box attacks: prompt nesting and multi-turn dialogue attacks. 1) Prompt Nesting Attack. Prompt nesting bypasses security features by nesting malicious intents in normal prompts, altering LLMs’ context. DeepInception Li et al. (2023b) exploit nested scenarios, while ReNeLLM Ding et al. (2024) rewrites prompt to jailbreak based on code completion, text continuation, or form-filling tasks. MJP Li et al. (2023a) uses multi-step approaches with contextual contamination to reduce moral constraints, prompting malicious responses. 2) Multi-turn Dialogue Attack. LLMs that are safe in isolated, single-round interactions can be gradually manipulated into generating harmful outputs through multiple rounds of interaction Russinovich et al. (2024); Zhou et al. (2024b); Yang et al. (2024); Cheng et al. (2024). Multi-turn dialogue attack leverages the multi-turn nature of conversational interactions to gradually erode an LLM’s content restrictions.
@@ 6 Conclusion
In this study, we enhance the understanding of jailbreak attacks by proposing a novel approach, AdVersArial meTAphoR (AVATAR) that manipulates LLMs to generate harmful content through calibrating benign metaphors, rather than generating harmful content from scratch. AVATAR consists of Adversarial Entity Mapping (AEM) and Metaphor-Induced Reasoning (MIR). Among them, AEM identifies metaphors by using the model crowdsourcing strategy, thus enabling the stable acquisition of suitable metaphors before jailbreaking the target LLM. MIR induces the target LLM to analyze metaphors, thereby generating harmful content. By further introducing the metaphor calibration mechanism, MIR improves the efficiency of transforming benign content into harmful forms. Experiments have demonstrated that our AVATAR is effective in generating transferable adversarial attacks on multiple advanced LLMs.
@@
While AVATAR achieves effective jailbreaks via metaphorical prompt manipulation without complex tricks or sophisticated jailbreak templates, several limitations remain:
Limited real-world grounding in metaphor mapping.
Since LLMs are trained on textual data alone, their representation of harmful domains may lack technical accuracy or practical relevance. To mitigate this, AVATAR incorporates a crowdsourcing strategy to increase the diversity and relevance of metaphor candidates.
@@
Built-in safety mechanisms may suppress key elements during toxic entity extraction, leading to incomplete or distorted representations. We address this by applying filtering techniques to remove alignment-related or non-informative terms during the mapping stage.